### v1

@@ Abstract
Inference on overall ranking of a set of entities, such as chess players, subpopulations or hospitals, is an important problem. Estimation of ranks based on point estimates of means does not account for the uncertainty in those estimates. Treating estimated ranks without regard for uncertainty is problematic. We propose a Bayesian solution. It is competitive with recent frequentist methods, and more effective and informative, and is as easy to implement as it is to compute the posterior means and variances of the entity means. Using credible sets, we created novel credible distributions for the rank vector of the entities. We evaluate the Bayesian procedure in terms of accuracy and stability in two applications and a simulation study. Frequentist approaches cannot take account of covariates, but the Bayesian method handles them easily.
@@ 1 Introduction
Reliable estimates of measures of important characteristics of agriculture, employment, health, education, for example, for various subpopulations are some of the important data produced by many national and international statistics offices around the world. For example, the United Nations Sustainable Development Goals, the US Small Area Income and Poverty Estimation (SAIPE) and many programs in European Union need accurate poverty data at disaggregated levels for making public policies and implementing welfare programs. Disaggregated levels may be defined by partitioning a particular population into several subpopulations. For example, subpopulations may refer to the states of a country (population). Effective implementations of many welfare programs and economic development policies critically need accurate point estimates of important characteristics at relevant subpopulation level. In this article we will also use “small areas” or “small domains” to refer to “subpopulations”. Definitely, the word “subpopulation” has a wider meaning. In a broader sense we use the word “entity” which may refer to a subpopulation or an individual. But in its place we will also use “small area” since our proposed solution to the problem of interest emerges from small area estimation methodology.
Productions of accurate estimates of the means for one or more characteristics for the subpopulations are usually the primary goals of many government agencies. However, inference on overall ranking of a set of entities, which may be individuals, small areas, hospitals, herds of animals, etc. is a problem of significant interest. Often it is crucially important to identify subpopulations with very low or high means of a characteristic. Effective implementation of welfare and well-being programs overseen by government or private organizations, identifying the better yielding crop varieties in agriculture research, or top-breeding sires in breeding research, etc. require reliable estimation of the rank vector. Due to limited resources government agencies may want to identify some of the most underprivileged or impoverished subpopulations to implement government programs to support those groups. Similarly, a business entity planning to open new locations would like to know the localities (or geographic regions) with the most demands for its products and services. In these cases, in addition to estimating the true subpopulation means, it is also important to accurately identify subpopulations that are either at the upper or at the lower end in terms of their means. This goal requires accurately estimating the ranks of several or all the subpopulations. According to Laird and Louis (1989), ranking can draw attention to unusually good or poor performance of some subpopulations. They noted that ranking of objects and entities arises in a wide variety of applications, and an appropriate solution provides an important tool to the investigators and policy makers setting priorities for the next stage of their investigations. The importance of joint ranking of $m$ subpopulations with means $\theta_{1},\cdots,\theta_{m}$ of a common characteristic has been emphasized in a pioneering article by Klein et al. (2020) (we will subsequently also use the abbreviation KWW to refer to this paper or the method considered by these authors). Overall true ranks of the subpopulations are unknown as those are determined by the true unknown $\theta_{i}$’s. Klein et al. (2020) noted that ranking of populations, governmental units, teams, players, health service providers, etc. based on experimental, observational or “survey data are usually released without direct statistical statements of uncertainty on estimated overall rankings”. They emphasized that prior to their work national statistical agencies did not have any statistical methodology to account for uncertainty in their released estimates of overall rankings. They further cautioned that from published point estimates of the means with no explicit ranking, practitioners frequently use the point estimates to naively ascertain ranks of the subpopulations. Ranks determined this way, which are only point estimates, ignore any uncertainty in the estimates of the true means $\theta_{i}$’s as these estimates are invariably subject to error, sampling error is the most common. Different sets of estimates of $\theta_{i}$’s from different data sets will potentially lead to different sets of ranks of the subpopulations. Clearly, even the best possible estimators of small area means are subject to error due to the sampling variability of the estimators. It is thus both imperative and a sound policy to evaluate uncertainty associated with the reported ranks based on reasonable estimates of subpopulation means. Obviously, the accuracy of the estimated ranks can be appropriately studied through the joint distribution of the ranks, determined from the estimators of the means. In an application to rank the fifty states of the U.S. and Washington, DC, using mean commuting times of workers sixteen or older and not working from home, Klein et al. (2020) used survey data collected from the American Community Survey (ACS). In their approach to the joint confidence set for the true rank vector, which they determine from an appropriate joint confidence set of the true means, the authors treated the unknown means as constants. They followed the classical approach to estimating the ranks where the true parameters $\theta_{i}$’s are considered to be fixed but unknown. Some of the early works on ranking and selection of population means are by Bechhofer (1954) and Gupta (1956) who adopted the frequentist approach (that is, the classical approach) to inference. While Klein et al. (2020) also used the frequentist approach for their application, they briefly mentioned in a remark that an analogous Bayesian version of their frequentist solution is possible. We have proposed here based on the Bayesian approach two novel and easily implementable solutions to this problem. Our proposed solutions enjoy better interpretation and based on some desirable measures (explained later) are at least as competitive as or better than the frequentist solution. The Bayesian approach treats the unknown subpopulation means as random and assigns them a suitable joint distribution. Following the Bayesian approach to ranking of several binomial populations by Bland and Bratcher (1968), Govindarajulu and Harvey (1974) presented a unified Bayesian treatment of the ranking and selection problem. Goel and Rubin (1977) proposed a Bayesian approach on selecting a subset containing the best subpopulation. Laird and Louis (1989) used an empirical Bayes (EB) approach to the ranking problem to environmental monitoring of regions or health service providers suspected of elevated disease rates for residents or failure rates from services for patients, requiring attention to mitigate risks. Indeed Morris and Christiansen (1996) highlighted the utility of Bayesian approach to the ranking of hospitals. Unlike the frequentist approach, the EB and the Bayesian approaches facilitate hierarchical modeling of the random latent subpopulation means through regression on relevant available covariates. An early application of hierarchical modeling to ranking of subpopulations is due to Efron and Morris (1975). In their application to rank thirty-six cities of El Salvador based on prevalence of toxoplasmosis rates, they indicated the pitfall of ranking the cities using only observed estimates of the rates without paying attention to the variability of these estimates. In the next section we provide a brief review of the relevant literature on ranking. We introduce two reasonable Bayesian models by specifying suitable joint prior distributions for the unknown subpopulation means. Corresponding to a prior distribution a Bayesian model is formed by augmenting the sampling distribution of the data, which is the cornerstone of the classical approach to a frequentist solution. In our Bayesian formulation, we consider two noninformative improper prior distributions. One prior distribution that provides no structure for the unknown $\theta_{i}$’s we consider it reasonable since joint credible sets for the $\theta_{i}$’s that are derived from the resulting posterior distribution exactly agree with the joint confidence sets of the $\theta_{i}$’s developed for the application by by Klein et al. (2020). The second prior distribution is based on a conjugate prior for the means leading to a widely used hierarchical model, introduced earlier by Fay and Herriot (1979), Berger (1985), Aitkin and Longford (1986), Berger and Deely (1988), Laird and Louis (1989) and Morris and Christiansen (1996). While most of these authors used an EB approach by estimating the parameters of the prior distribution in the hierarchical model, Berger and Deely (1988) and Morris and Christiansen (1996) used the hierarchical Bayes (HB) approach (or its approximation) by assigning reasonable noninformative improper priors to the model parameters. We follow these latter groups and will pursue a fully Bayesian (that is, HB) approach with similar improper priors on the model parameters. We add that whereas Berger and Deely (1988) and Morris and Christiansen (1996) used the HB approach for (point) estimation of individual ranks, we consider this approach for joint credible set of the overall mean vector by proposing a credible distribution through an innovative idea.
@@ 2 A selected review of the ranking literature and introduction of Bayesian models
Our motivation for this research has stemmed from a need to developing a Bayesian solution to the joint ranking problem proposed in a pioneering paper by Klein et al. (2020). A population may be partitioned into $m$ subpopulations. Suppose $\theta_{i}$ is the mean for certain characteristic of the $i$th subpopulation. In the application considered by Klein et al. (2020), they are interested in ranking 51 subpopulations, the fifty states of US and Washington, DC, based on $\theta_{i}$, the mean commuting times of the workers not working from home, $i=1,\cdots,m$. Similar to theirs, our goal is to make joint inference about the true ranks $\check{r}_{1},\cdots,\check{r}_{m}$, of the $m$ subpopulations based on estimators $Y_{1},\cdots,Y_{m}$ of $\theta_{1},\cdots,\theta_{m}$, and appropriate models. Alternatively, these $m$ subpopulations may refer to $m$ individuals. In ranking $\theta_{i}$’s we treat them continuous variables. We rank the $\theta_{i}$’s from the smallest to the largest, where the rank of $\theta_{i}$ is
@@
The estimated overall ranking based on the estimators $Y_{1},\cdots,Y_{m}$ of $\theta_{1},\cdots,\theta_{m}$, is denoted by $R_{1},\cdots,R_{m}$, where
@@
We already mentioned that the estimators $R_{i}$’s are subject to errors, which stem from the errors in the point estimators $Y_{i}$’s. It is important to account for uncertainty in the estimators $R_{i}$’s. Klein et al. (2020) proposed a clever solution to account for uncertainty in the joint estimation of the true rank vector. While they adopted a frequentist approach, we will present purely Bayesian solutions as alternatives based on two Bayesian models that treat the unknown $\theta_{i}$’s random. Actually, to estimate individual ranks $\check{r}_{i}$’s of the $\theta_{i}$’s, Aitkin and Longford (1986) considered setups where in one case they assumed $\theta_{i}$’s are constants, and in the second case they modeled $\theta_{i}$’s by some probability distribution, possibly depending on parameters ($\psi$, say). The latter case constitutes a hierarchical modeling for the $Y_{i}$’s and $\theta_{i}$’s, and estimates of $\theta_{i}$’s or $\check{r}_{i}$’s can be determined by an EB approach if $\psi$ is estimated, or a fully Bayesian (FB) approach if $\psi$ is also assigned a prior distribution, proper or improper. Aitkin and Longford (1986) followed the EB approach for hierarchical model. Berger and Deely (1988) proposed a fully Bayesian approach to test $H_{0}$ for equality of subpopulation means, and conditional on $H_{0}$ is false, the authors estimated probability of $\theta_{i}$ being the largest among the means, for $i=1,\cdots,m$. They estimate the ranks of the $\theta_{i}$’s based on the ranks of these probabilities. In the remainder of the article, for compactness of notation, we will often use $Y,y,\theta$ and $\check{r}$ to denote the vectors $(Y_{1},\cdots,Y_{m})^{T},(y_{1},\cdots,y_{m})^{T},(\theta_{1},\cdots,\theta_{m})^{T}$ and $(\check{r}_{1},\cdots,\check{r}_{m})^{T}$, respectively. For the unknown true means $\theta_{1},\cdots,\theta_{m}$, with corresponding estimators $Y_{1},\cdots,Y_{m}$, Klein et al. (2020) assumed that
@@
where the variances $D_{i}$ associated with $Y_{i}$ are assumed to be known. In the small area estimation literature $D_{i}$’s are often referred to as the sampling variances. To account for uncertainty of the rank estimators $R_{1},\cdots,R_{m}$, the authors constructed, for some specified $\alpha\in(0,1)$, a $(1-\alpha)$-level joint confidence set for $\theta$ based on confidence intervals $I_{i}=(Y_{i}-z_{1-\frac{\gamma}{2}}\sqrt{D_{i}},Y_{i}-z_{\frac{\gamma}{2}}\sqrt{D_{i}})\equiv(L_{i},U_{i}),i=1,\cdots,m,$ for individual $\theta_{i}$’s, where $z_{q}$ is the $q$th quantile of the standard normal distribution. These authors determined $\gamma$ from $\alpha$ in two ways: by applying (i) Bonferroni’s inequality, and (ii) independence of the $Y_{i}$’s, so that the joint confidence level is $(1-\alpha)$. Using the individual intervals, they formed the Cartesian product $I_{1}\times I_{2}\cdots\times I_{m}$ to obtain the joint confidence set of the true mean vector $\theta$. They used this confidence set to develop a $(1-\alpha)$-level joint confidence set for the true rank vector $\check{r}$. We briefly review their method in Section 3 below. We note that Bonferroni’s method does not need independence of the pivotal quantities to create intervals for the individual means $\theta_{i}$’s, it is broadly applicable. However, the method generates intervals that are usually slightly wider than those produced by the independent method. This may result in additional overlapping intervals and a slightly more expansive set for overall ranking (cf. the rank sets for these two methods for HI and TX, in Table 2 of KWW). While the frequentist joint confidence set of the rank vector developed by Klein et al. (2020) accounts for the uncertainty in the estimated rank vector $(R_{1},\cdots,R_{m})$, some limitations to their method are that (i) it does not provide any opportunity to utilize any relevant covariate that may be related to the $\theta_{i}$’s, and (ii) the generated set is overly “stretched out”. As mentioned earlier, following Berger and Deely (1988) we will assign to the $\theta_{i}$’s a similar probability model and present a fully Bayesian solution for a joint credible set estimation of the overall ranking. We emphasize that although Aitkin and Longford (1986), Berger and Deely (1988), Laird and Louis (1989) and Morris and Christiansen (1996) considered EB/HB estimation of individual ranks based on hierarchical modeling, none of them considered the (credible) set estimation of the rank vector $(\check{r}_{1},\cdots,\check{r}_{m})^{T}$. Our proposed solution, explained later, is a credible distribution.
2.1 An unstructured Bayesian model
Specification of the unstructured Bayesian model is simple. We augment the sampling model in equation (3) by the prior pdf for $\theta_{1},\cdots,\theta_{m}$

### v2

@@ Abstract
Inference on overall ranking of a set of entities, such as athletes or players, schools and universities, hospitals, cities, restaurants, movies or books, companies, states, countries or subpopulations, based on appropriate characteristics or performances, is an important problem. Estimation of ranks based on point estimates of means does not account for the uncertainty in those estimates. Treating estimated ranks without any regard for uncertainty is problematic. We propose a novel solution using the Bayesian approach. It is easily implementable, competitive with a popular frequentist method, more effective and informative. Using suitable joint credible sets for entity means, we appropriately create credible distributions (CDs, a phrase we coin), which are probability distributions, for the rank vector of entities. As a byproduct, the supports of the CDs are credible sets for overall ranking. We evaluate our proposed procedure in terms of accuracy and stability using a number of applications and a simulation study. While the frequentist approach cannot utilize covariates, the proposed method handles them routinely to its benefit.
@@ 1 Introduction
Reliable estimates of measures of important characteristics of agriculture, employment, health, education, for example, for various subpopulations are some of the important data produced by many national and international statistics offices around the world. For instance, the United Nations Sustainable Development Goals, the US Small Area Income and Poverty Estimation (SAIPE) and many programs in European Union need accurate poverty data at disaggregated levels for making public policies and implementing welfare programs. Disaggregated levels may be defined by partitioning a particular population into several subpopulations. For example, subpopulations may refer to the states of a country (population). While the term “subpopulations” has a wider meaning, we will also use “small areas” or “small domains” in its place. In a broader sense we use the word “entity” which may refer to a subpopulation or an individual.
Accurately estimating the means for one or more characteristics for the subpopulations is usually the primary goal of the government agencies. However, inference on overall ranking of entities, such as individuals, small areas, hospitals, herds of animals, etc. is also substantially important. Often it is also important to identify entities with very low or high means. Effective implementation of welfare or well-being programs overseen by authorities, identification of better yielding crop varieties in agriculture, or the top-breeding sires in animal breeding, etc. require reliable estimation of the rank vector. Limited resources mandate authorities to identify the underprivileged or impoverished subpopulations to provide assistance. Similarly, a business to open new stores/offices would like to know the localities with the most demands for its products and services.
These goals require accurately estimating the ranks of many or all subpopulations. According to Laird and Louis (1989), ranking draws attention to unusually good or poor performance of some subpopulations. They noted that ranking arises in a wide variety of applications, and appropriate solutions provide useful tools to investigators and policy makers for setting priorities for the next stage of their investigations.
The importance of joint ranking of $m$ subpopulations with means $\theta_{1},\cdots,\theta_{m}$ of a common characteristic has been emphasized in a pioneering article by Klein et al. (2020) (we will subsequently also use the abbreviation KWW to refer to this paper or the method by these authors). Overall true ranks of the subpopulations are unknown as those are determined by the true unknown $\theta_{i}$’s. Klein et al. (2020) noted that ranking of populations, governmental units, teams, players, health service providers, etc. based on experimental, observational or “survey data are usually released without direct statistical statements of uncertainty on estimated overall rankings”. They emphasized that prior to their work national statistical agencies did not adopt any statistical methodology to account for uncertainty of released estimates of overall rankings. They further cautioned that from published point estimates of the means only, practitioners frequently use those to naively rank the subpopulations. Such ranks are only point estimates, and ignore uncertainty in the estimates of $\theta_{i}$’s. Different sets of estimates of $\theta_{i}$’s from different data sets will potentially lead to different sets of ranks of the subpopulations. Clearly, even the best possible estimators of small area means are subject to error due to sampling variability of the estimators. It is thus both imperative and a sound policy to evaluate uncertainty associated with the reported ranks based on reasonable estimates of subpopulation means. Obviously, the accuracy of the estimated ranks can be appropriately studied through the joint distribution of the ranks, induced by the joint distribution of estimators of the means.
In an application to rank the fifty states of the U.S. and Washington, DC, using mean commuting times of workers who are sixteen or older and not working from home, Klein et al. (2020) used data from the American Community Survey (ACS). In their approach to the joint confidence set for the true rank vector, which they determine from an appropriate joint confidence set of the true means, the authors treated the unknown means non-random. They followed the classical approach where the unknown true parameters $\theta_{i}$’s are non-random. Some of the early frequentist works on ranking and selection of population means are by Bechhofer (1954) and Gupta (1956). While Klein et al. (2020) first created a joint confidence set for the mean vector $\theta=(\theta_{1},\cdots,\theta_{m})^{T}$ to create confidence set for ranks based on the frequentist approach, they briefly remarked that an analogous Bayesian version of their frequentist solution is possible through construction of a joint credible set for $\theta$. Our Bayesian solution is not based on what Klein et al. (2020) alluded to in that remark. Our solution is novel and to implement it we need a Bayesian paradigm. We have proposed here two easily implementable solutions to this problem.
In our Bayesian approach, we first obtain an empirical credible set for $\theta$ by selecting a subset of $\theta$ values from an MCMC sample that are in a suitable region. Subsequently, we use these selected values to create probability distributions of the components of the rank vector. To convey these probability distributions originate from a credible region we coined the phrase “credible distributions” (CDs) to refer to these distributions. Our proposed solutions admit better interpretations and appear to be more accurate based on several desirable measures (see Sections 6 to 10). They are at least as good as, if not better than, the frequentist solution. A reviewer brought it to our attention to a very recent paper by Mogstad et al. (2024) on frequentist approach to overall ranking. These authors extended the KWW work by considering joint confidence intervals of all pairwise differences of the form $\theta_{i}-\theta_{j}$. As in the approach by KWW, the confidence intervals of the differences of the means are used to create joint confidence estimation of overall ranking. Since our objective is to develop an intuitive and easily implementable Bayesian solution to the overall ranking, we stick here only to the seminal work by Klein et al. (2020) for comparison with our Bayesian solution.
The Bayesian approach treats $\theta$ as random and assigns them a suitable joint distribution. Following the Bayesian approach to ranking several binomial populations by Bland and Bratcher (1968), Govindarajulu and Harvey (1974) presented a unified Bayesian treatment of the ranking and selection problem. Goel and Rubin (1977) proposed a Bayesian approach on selecting a subset containing the best subpopulation. Laird and Louis (1989) used an empirical Bayes (EB) approach to the ranking problem to environmental monitoring of regions suspected of elevated disease rates for residents or health service providers suspected of elevated services failure rates for patients, requiring attention to mitigate risks. Indeed, Morris and Christiansen (1996) highlighted the utility of Bayesian approach to the ranking of hospitals. Unlike the frequentist approach of Klein et al. (2020), the EB and the (fully) Bayesian approaches facilitate hierarchical modeling of the random latent means $\theta$ through regression on relevant available covariates. An early application of hierarchical modeling to ranking of subpopulations is due to Efron and Morris (1975). In their application to rank thirty-six cities of El Salvador based on prevalence of toxoplasmosis rates, they indicated the pitfall of ranking using only observed estimates of the rates but ignoring the variability of these estimates.
In Section 2 we provide a brief review of the relevant literature on ranking. We introduce two reasonable Bayesian models by specifying suitable joint prior distributions for the unknown subpopulation means. Corresponding to a prior distribution a Bayesian model is formed by augmenting the sampling distribution of the data, which is the cornerstone of the classical approach to a frequentist solution. In our Bayesian formulation, we consider two noninformative improper prior distributions. The prior distribution in equation (4) that provides no structure for the unknown $\theta_{i}$’s is justified since joint credible sets for the $\theta_{i}$’s from the posterior distribution exactly agree with the joint confidence sets of the $\theta_{i}$’s developed by Klein et al. (2020). The second prior distribution in Section 2.2 is based on a conjugate prior for the means that leads to a widely used hierarchical model, introduced earlier by Fay and Herriot (1979), Berger (1985), Aitkin and Longford (1986), Berger and Deely (1988), Laird and Louis (1989) and Morris and Christiansen (1996). While most of these authors used an EB approach by estimating the parameters of the prior distribution in the hierarchical model, Berger and Deely (1988) and Morris and Christiansen (1996) used the hierarchical Bayes (HB) approach (or its approximation) by assigning reasonable noninformative, minimally influential, improper priors to the model parameters.
We follow these latter groups and pursue a fully Bayesian (that is, HB) approach with similar, minimally influential improper priors on the model parameters. We add that whereas Berger and Deely (1988) and Morris and Christiansen (1996) used the HB approach for (point) estimation of individual ranks, we consider this approach for joint set estimation of the rank vector. Our solution, elaborated in Section 4, is based on empirical representation of a $100(1-\alpha)\%$ credible set for $\theta$ using MCMC samples of $\theta$ from the posterior distribution. The MCMC samples that are in the credible region are taken to represent the empirical credible region (ECR). Each of these samples in the ECR is ranked, and in totality they induce relative frequency distributions, defined earlier as credible distributions, of the ranks of the entities.
Here is a brief outline of the paper. In Section 2 we introduce some useful notation. In Subsections 2.1 and 2.2 we present two Bayesian models that are the building blocks of our solution. In Section 3 we present a brief review of the Klein et al. (2020) frequentist solution. In Section 4 we describe empirical credible sets. In Subsection 4.1 we discuss empirical Cartesian and elliptical (related to highest posterior density) credible sets for $\theta$. In Section 5 we describe computation of credible distributions (CDs). In Section 6 we compare various set estimators (confidence and credible) of $\theta$ based on suitable size measures. In Section 7 we introduce two new measures to compare the frequentist and Bayesian solutions to the overall ranking problem. In Section 8 we apply and contrast our solution to the Klein et al. (2020) solution for ranking of the US states based on commuting times. In Section 9 we develop credible distributions to rank the U.S. states based on median incomes. In Section 10 we develop credible distributions for the baseball data studied by Efron and Morris (1975). In Subsection 10.1 we present simulation results under the baseball example setup to compare various solutions for overall ranking. Section 11 provides a summary and a few concluding remarks. In Subsection S.2 we review some results on volume and derive average representative length of an elliptical set. In Subsection S.4 we present a full discussion on the baseball example.
@@ 2 A selected review of the ranking literature and introduction of Bayesian models
Our motivation for this research is to develop a Bayesian solution to the joint ranking problem proposed in a pioneering paper by Klein et al. (2020). Suppose $\theta_{i}$ is the mean for certain characteristic of the $i$th subpopulation, $i=1,\cdots,m$. In their application, Klein et al. (2020) are interested in ranking the fifty states of US and Washington, DC, based on $\theta_{i}$, the mean commuting times of the workers not working from home, $i=1,\cdots,m=51$. Similar to theirs, our goal is also to make joint inference about the true ranks $\check{r}_{1},\cdots,\check{r}_{m}$ (see equation (1)), of the $m$ subpopulations based on estimators $Y_{1},\cdots,Y_{m}$ of $\theta_{1},\cdots,\theta_{m}$, and appropriate models. Alternatively, these subpopulations may be called entities. In ranking $\theta_{i}$’s we treat them continuous variables. We rank the $\theta_{i}$’s from the smallest to the largest, where the rank of $\theta_{i}$ is
@@
Based on $Y_{1},\cdots,Y_{m}$, estimators of $\check{r}_{1},\cdots,\check{r}_{m}$ are $R_{1},\cdots,R_{m}$, where
@@
Obviously, the estimators $R_{i}$’s are subject to random errors, which stem from the point estimators $Y_{i}$’s. It is important to account for uncertainty in the estimators $R_{i}$’s. Klein et al. (2020) proposed a clever solution to account for uncertainty in the joint estimation of the true rank vector. While they adopted a frequentist approach, we will present purely Bayesian solutions as alternatives based on two Bayesian models for the $\theta_{i}$’s. Actually, to estimate individual ranks $\check{r}_{i}$’s of the $\theta_{i}$’s, Aitkin and Longford (1986) considered setups where in one case they assumed $\theta_{i}$’s are non-random, and in the second case they modeled $\theta_{i}$’s by some probability distribution, possibly depending on parameters ($\psi$, say). The latter case constitutes a hierarchical modeling for the $Y_{i}$’s and $\theta_{i}$’s, and estimates of $\theta_{i}$’s or $\check{r}_{i}$’s can be determined by an EB approach if $\psi$ is estimated, or a fully Bayesian (FB) approach if $\psi$ is also assigned a prior distribution, proper or improper. Aitkin and Longford (1986) followed the EB approach for hierarchical model. Berger and Deely (1988) proposed a fully Bayesian approach to test a null hypothesis ($H_{0}$) for equality of subpopulation means, and conditional on $H_{0}$ is false, they estimated probability of $\theta_{i}$ being the largest of the $m$. They estimate the ranks of the $\theta_{i}$’s based on the ranks of these probabilities.
In the remainder of the article, for compactness of notation, we will often use $Y,y,\theta$ and $\check{r}$ to denote the vectors $(Y_{1},\cdots,Y_{m})^{T},(y_{1},\cdots,y_{m})^{T},(\theta_{1},\cdots,\theta_{m})^{T}$ and $(\check{r}_{1},\cdots,\check{r}_{m})^{T}$, respectively. Here, $y$ is the observed value of the random vector $Y$. Klein et al. (2020) assumed for the estimators $Y_{1},\cdots,Y_{m}$, that
@@
where the variances $D_{i}$ associated with $Y_{i}$ are assumed to be known. In the small area estimation literature $D_{i}$’s are often referred to as the sampling variances.
To account for uncertainty of the rank estimators $R_{1},\cdots,R_{m}$, the authors first constructed, for some specified $\alpha\in(0,1)$, a $(1-\alpha)$-level joint confidence set for $\theta$ based on confidence intervals $I_{i}=[Y_{i}-z_{1-\frac{\gamma}{2}}\sqrt{D_{i}},Y_{i}-z_{\frac{\gamma}{2}}\sqrt{D_{i}}]\equiv[L_{i},U_{i}],i=1,\cdots,m,$ for $\theta_{i}$’s, where $z_{q}$ is the $q$th quantile of the standard normal distribution. They determined $\gamma$ from $\alpha$ in two ways: by applying (i) Bonferroni’s inequality ($\gamma=\alpha/m$), and (ii) independence of the $Y_{i}$’s, so that the joint confidence level is $(1-\alpha)(=(1-\gamma)^{m})$. Using the individual intervals, they formed the Cartesian product $I_{1}\times I_{2}\cdots\times I_{m}$ to obtain the joint confidence set of the true mean vector $\theta$. They used this confidence set to develop a $(1-\alpha)$-level joint confidence set for the true rank vector $\check{r}$. We briefly review their method in Section 3 below. Bonferroni’s method does not need independence of the pivotal quantities to create confidence set for $\theta$ and it is broadly applicable. However, the intervals are usually slightly wider than those produced by the independent method. This may result in additional overlapping intervals and a slightly more expansive set for overall ranking (cf. the rank sets for these two methods for HI and TX, in Table 2 of KWW).
While the frequentist joint confidence set developed by Klein et al. (2020) accounts for the uncertainty in the estimated rank vector $(R_{1},\cdots,R_{m})$, some limitations to their method are that (i) it does not provide any opportunity to utilize any relevant covariate that may be related to the $\theta_{i}$’s, and (ii) the generated set is overly “stretched out”. As mentioned earlier, following Berger and Deely (1988) we will assign to the $\theta_{i}$’s a similar probability model and present a fully Bayesian solution for a joint set estimation of the overall ranking. We emphasize that although Aitkin and Longford (1986), Berger and Deely (1988), Laird and Louis (1989) and Morris and Christiansen (1996) considered EB/HB estimation of individual ranks based on hierarchical modeling, none considered the (credible) set estimation of the rank vector $(\check{r}_{1},\cdots,\check{r}_{m})^{T}$. Our proposed solution, defined and explained in Section 5, is based on credible distributions.
2.1 An unstructured Bayesian (UB) model
Here, we augment the sampling model in equation (3) by the prior pdf for $\theta_{1},\cdots,\theta_{m}$